OpenStudy (anonymous):

Sabrina goes into a bank and gets change for $50. She gets the change in 5's and 1's and there are 22 bills total. How many 5's are there and how many 1's are there. The thing is, I know that answer, but I don't know how to show the work. :-/ Please help me to understand how to show the work. There are 15- 1 dollar bills and 7- 5 dollar bills

OpenStudy (compassionate):

5x + 1x = 22 Is the equation

OpenStudy (rajee_sam):

First you need to set up the system of equations

OpenStudy (anonymous):

after I get the equation, I'm lost. I just did the process of multiplication/ division with addition and subtraction to get the answer. But as far as getting any further than that 5x+1x=22.... I'm lost

OpenStudy (rajee_sam):

let the no. of 5 dollor bills be X and the no. of 1 Dollar Bills be Y. Those are the unknowns you need to find. It is given X + Y = 22 Also we know 5X + 1Y = 50 Now you have two equations and you have to solve

OpenStudy (rajee_sam):

Do you know how to do that?

OpenStudy (anonymous):

Pretty much, when I started the equation and got the 5x+1x written down, I went 10 ones plus 8 fives is 18 bills but still 50 dollars. Took it up to 15 ones and down to 7 fives and got the answer. Anything between all that, I'm not sure how to do or what to do.

OpenStudy (compassionate):

x + y = 22 5x + 1y = 50 Solve using elimination

OpenStudy (rajee_sam):

If you do trial and error it is going to take you a whole lot of time and thats where algebra helps you. Now did you understand how I set up the equations?

OpenStudy (anonymous):

Thanks compassionate. Sam, I understand how to set everything up... showing how I got to the answer (the way this teacher wants) is a different matter and is what I'm needing help figuring out. I could show him how I did it all day long but it wouldn't be right. I got to the equations, but couldn't figure out what to do next.

OpenStudy (rajee_sam):



OpenStudy (anonymous):

Wow! Thank you! That helps a lot.